OpenStudy (anonymous):

give the slope intercept form of the equation of the line that is perpendicular to 7x - 3y - 18 and contains F(6, 8)

OpenStudy (anonymous):

Rearrange 7x-3y-18 in the form y=mx+b Know how to do that?

OpenStudy (anonymous):

literally i struggle so badly with math so no help me please

OpenStudy (anonymous):

Okay step by step 1) Rearranging equation to look like y=mx+b: 7x-3y-18=0 --> 7x-18=3y ---> 3y=7x-18 ---> y= (7/3)x -6 1) is y=7/3x-6 2) Rule to remember: to find a line perpendicular to another it is the slope flipped (the reciprocal), multiplied by (-1) in this case the slope of the first line is 7/3. So the perpendicular line would be -3/7 So the line perpendicular has a line equation of y=-3/7x +b 3) find b plug x 6 and y 8 in the equation y=-3/7x + b and find b 8=(-3/7)*6+ b --> 8=-2.57+b ---> b=10.57 Perpendicular line equation is y=-3/7x+10.57

OpenStudy (texaschic101):

7x - 3y = 18 -3y = -7x + 18 (divide by -3) y = -7/-3x + 18/-3 y = 7/3x - 6 the slope of this equation is 7/3, however, since we are looking for a perpendicular line, we need to get the negative reciprocal of that slope. How we do that is " flip " the slope and change the sign. So we have slope 7/3 and we need the negative reciprocal slope which is -3/7 (see how I flipped the slope and changed the sign.) Now we will use point slope formula... y - y1 = m(x - x1) using slope -3/7 (6,8) x1 = 6 and y1= 8 now lets sub y - 8 = -3/7(x - 6) y - 8 = -3/7x + 18/7 y = -3/7x + 18/7 + 8 y = -3/7x + 18/7 + 56/7 y = -3/7x + 74/7

OpenStudy (anonymous):

that fraction seems wrong are you sure?

OpenStudy (anonymous):

Same as what i got, but I put in decimal form.

OpenStudy (texaschic101):

unless you want to turn it into a mixed number....74/7 = 10 4/7

OpenStudy (anonymous):

would they both be correct than

OpenStudy (texaschic101):

yes

OpenStudy (texaschic101):

you can put it in decimal, mixed number, or improper fraction

OpenStudy (anonymous):

@texaschic101 are you good with stats? I could use some help :/

OpenStudy (anonymous):

youre a saint think you could help me with some others?

OpenStudy (anonymous):

I gotta focus on my own work now, or I would! :(

OpenStudy (texaschic101):

I am not so good at stats...sorry

OpenStudy (anonymous):

Thanks anyways.